# Far from Perfect: Quantum Error Correction with (Hyperinvariant) Evenbly Codes

### v1

@@ III.3 Evenbly Codes as Subsystem Codes
[…]
@@
We show below that this constant-rate version of the Evenbly code yields a rate of $R=k/n=1/8$; for comparison, the trivial gauge subsystem variant (i.e. the maximum number of bulk logical indices are used for storing logical information and gauge group is empty) of the $\{5,4\}$ Evenbly code exhibits an asymptotic rate of $R=0.866$. Such finite encoding rates are a common feature of codes defined on hyperbolic manifolds [54, 8]. Both of the code variants’ threshold properties will be evaluated in Section III.6. In both cases, $\mathcal{H}_{\text{bulk}}$ refers to all other logical qubits partitioned into the gauge subsystem.
Figure 6: Gauged Evenbly codes considered in this paper. (a) A zero-rate Evenbly code where only the central logical qubit (red dot) is considered, all other bulk legs being gauge-fixed (gray disks). (b) A constant-rate Evenbly code, where each logical qubit occupies a unique crossing point in the bulk. Here, hyperbolic geodesics only pass over sites that are gauge-fixed. An example for such a pair of geodesics for one logical qubit is shaded in blue. (c) In the $X$ gauge, all the gauge-fixed tensors act as pairs of overlapping EPR pairs, mapping each logical qubit to four unique physical boundary sites under the initial $\llbracket 4,1,2\rrbracket$ code. Any other gauge mixes the encodings of logical qubits. (d) A variant of the code in (b), with half of all logical qubits on every layer gauge-fixed. We examine this code’s threshold under erasure in Section III.6.
@@
III.4 Rate and Distance Scaling
Calculating the asymptotic rate of a $\{5,4\}$ Evenbly code can be done by following the prescriptions found in [14, 8, 45], and is related to an inflation procedure for hyperbolic tilings known as vertex inflation [44, 45]. As in these previous examples, rates can be evaluated by assessing the tensorial content in the bulk, and comparing with the total number of boundary sites.
In each layer of the tiling, we can distinguish vertex tensors (and their logical qubits) by the number of legs going out towards the next layer (for the last layer, these open legs support the physical qubits). We denote vertices with one to four outgoing legs with $\alpha,\beta,\gamma$ and $\delta$. For vertex inflation, $\delta$ only appears as the central (seed) tensor and $\gamma$ never, following the substitution rules
@@ III.4 Rate and Distance Scaling
Vertex inflation determines a substitution matrix $M$; every column of $M$ represents the number of sites of each type that arise in every subsequent layer as a result of applying the inflation rules. Combining this matrix with the vector $\eta=(n_{\alpha},n_{\beta},n_{\delta})^{\text{T}}$ which represents the number of sites at the $L^{\text{th}}$ layer, we construct a recursion relation of the form
$\eta^{(L+1)}\equiv\begin{pmatrix}n_{\alpha}^{(L+1)}\\ n_{\beta}^{(L+1)}\\ n_{\delta}^{(L+1)}\end{pmatrix}=\begin{pmatrix}1&2&4\\ 1&3&8\\ 0&0&0\end{pmatrix}\begin{pmatrix}n_{\alpha}^{(L)}\\ n_{\beta}^{(L)}\\ n_{\delta}^{(L)}\end{pmatrix}=M\eta^{(L)}\ .$ (10)
@@
This is the rate of the max-rate Evenbly code in which all logical qubits are gauge-free. However, in Section III.6, we treat another variant, where a number of logical qubits are gauge-fixed. The remaining logical qubits correspond to every second $\beta$-type vertex in the language of inflation steps. Confirming the geometric argument of Figure 6(b), we find that this code has the constant rate
@@
However, since every layer is half-filled (Figure 6(d)), we must then divide by two, leaving the rate $R^{\prime}=1/8$.
We now consider code distances, i.e., the minimum physical weight of any logical operator, distinguishing between bit distances of single-qubit logical operators and word distances of arbitrary multi-qubit logical operators. For the Evenbly code, distances heavily depend on the choice of gauge: In the max-rate case, acting with two physical Pauli operators can create a logical operation arbitrarily deep in the bulk, leading to $d_{\text{word}}=2$. If the two Pauli operators act on neighboring physical qubits, they will affect only logical qubits adjacent to the boundary (e.g. $Z_{j}Z_{j+1}$ acting as a $\bar{Z}$ on a single tile); however, if they act further apart, the result is a string of logical operators along a shortest path throughout the bulk, a feature that is related to the appearance of non-trivial 2-point functions in such codes [24].
@@
$w^{(L)}(\bar{X})$ $=(1+\sqrt{2})^{L}+(1-\sqrt{2})^{L}\ ,$ (22)
$w^{(L)}(\bar{Z})$ $=(1+\sqrt{2})^{L+1}+(1-\sqrt{2})^{L+1}\ .$ (23)
@@
$d_{\text{bit}}\sim(1+\sqrt{2})^{L}$ $=\left(\frac{n}{2+2\sqrt{3}}\right)^{\log_{2+\sqrt{3}}(1+\sqrt{2})}$ (24)
@@
The distance for $\bar{Z}$ scales similarly, with an additional prefactor of $1+\sqrt{2}$. To achieve a more useful code, we would like to increase $d_{\text{word}}$ for a greater resilience against errors while decreasing $d_{\text{bit}}$ so that targeted logical operations are easier to perform.
This condition is exactly what gauge-fixing some logical qubits of the Evenbly code achieves: In the most extreme example of the zero-rate code with only a single gauge-free logical qubit in the center of the tiling, $d_{\text{word}}=d_{\text{bit}}$ while being upper-bounded by the max-rate result from Equation 24. This upper bound may not be tight, as the availability of additional stabilizer terms on gauge-fixed tiles allows for more compact representations of logical operators on the boundary. Exactly computing these distances is a computationally difficult problem, though we expect that gauge-fixing in a Z or $Y$ gauge will reduce the power coefficient in Equation 24 somewhat below $0.67$.
@@
III.5 Erasure Threshold for a Zero-Rate Code
We analyze the threshold performance of the $\{5,4\}$ Evenbly code using two different decoding strategies for erasure errors: A greedy reconstruction method generalizing the approach of [8], in which the bulk logical qubits are iteratively reconstructed from the boundary and whose threshold we denote by $p^{\text{greedy}}_{\text{erasure}}$, and a Gaussian elimination algorithm adapted from [9] with threshold $p^{\text{gaussian}}_{\text{erasure}}$. The implementation and simulation details for both algorithms can be found in Appendix E. To summarize, greedy reconstruction relies on simple geometric steps but is less effective than the Gaussian elimination approach, which considers all possible physical representations of logical operators.
@@ III.5 Erasure Threshold for a Zero-Rate Code
Simulation results for recovery rates with both decoding algorithms under an erasure channel are shown in Figure 7, considering only the recovery of the central logical qubit. In subfigures (a)-(d), the threshold results $p_{\text{erasure}}$ are shown for each choice of gauge fixing discussed in Section III.3. No gauge fixing (subfigure (a)) corresponds to the max-rate case, whereas (b)-(d) consider different gauges of the zero-rate code. In agreement with the discussion of a ququart Evenbly code in [24], the max-rate case is highly susceptible to erasure errors and does not display a threshold.
@@
The logical $Y$ and $Z$ gauges (subfigures (c)-(d)) are different: Here we find a threshold close to the maximal value of $50\%$ for full reconstruction, with the simpler greedy reconstruction only achieving around $24\%$ and $40\%$, respectively. This relative underperformance of the greedy algorithm is due to its reconstruction of the bulk algebra on edges tile-by-tile, which may be prevented by erasure errors even though the logical operators of the central qubit are still recoverable. For this reason, the greedy recovery probability always lower-bounds full recovery, as already seen in other work [14, 9]. Conversely, the Gaussian elimination algorithm poses an upper bound to the recovery probability, as it has been shown to be optimal for erasure noise [9, 27].
III.6 Erasure Threshold for a Constant-Rate Code
As shown in Figure 6(b), one may also gauge-fix some of the logical qubits while keeping the rate of the code finite, leading to the constant-rate setting described in Section III.3. Keeping the two graph geodesics across any logical qubit gauge-fixed on all other bulk sites, we potentially increase the protection of each logical qubit from erasures closest to it on the boundary. In the $X$ gauge, this behavior directly generalizes the zero-rate case: Again, the gauged sites turn into EPR pairs that directly map a $\llbracket 4,1,2\rrbracket$ code onto four unique physical qubits, but with $k=n/4$ logical qubits instead of one. This is visualized in Figure 6(c).
[…]
@@ III.6 Erasure Threshold for a Constant-Rate Code
III.7 Threshold for Depolarizing and Pure Pauli Noise
Figure 9: Recovery curves for the zero-rate $\{5,4\}$ Evenbly code for depolarizing and biased Pauli noise. (a) A threshold appears at approximately $19.1\pm 0.94\%$ for depolarizing noise. (b) For pure Pauli $X$ noise, a preliminary threshold appears at approximately $50\%$. The same behavior was observed for pure Pauli $Y$ and $Z$ noise as well.
@@ III.7 Threshold for Depolarizing and Pure Pauli Noise
In Figure 9, (a) displays the recovery threshold $p_{\text{depo}}$ obtained as a result of varying the physical noise parameter. As is shown, a threshold appears in the region of approximately $p_{\text{depo}}\approx 19.1\pm 0.94\%$; this threshold value is very competitive with known code-capacity thresholds for the surface code, color code, and the holographic $\llbracket 6,1,3\rrbracket$ code [6, 13, 30]. Subfigures (b-d) display the results obtained for pure Pauli noise using the integer-optimization decoder. As can be ascertained, preliminary estimates indicate pure Pauli noise thresholds of approximately $50\%$, in line with the results obtained for the erasure channel. However, as we were limited in the computational power and time available, we report these findings here as preliminary, with three crossing points. However, if confirmed, it would seem that the Evenbly code is capable of reaching and exceeding known achievable bounds for random zero-rate quantum codes (known as the zero-rate hashing bound [31, 30]).
IV Discussion
IV.1 qCFTs & Holography
@@ IV Discussion
The construction presented in this work is one among many possible realizations of HTN codes; indeed, it was shown in [26] that many possible tensor decompositions for the original HTN ansatz exist, even those which are non-unitary in terms of individual tensor properties. However, constructing a quantum error-correcting code with non-unitary tensors may prove to be very challenging, and even more so when trying to determine relevant error correction properties of the resulting code.
@@
Although we have realized an exact holographic code which exhibits genuine polynomially-decaying boundary-to-boundary correlation functions, we have yet to investigate whether CFT data relevant for a known minimal-model CFT [47], or a quasiperiodically disordered qCFT [46], can appear as a result of the gauge-fixing techniques described. We leave this issue for future work.
@@
Evenbly codes at their core circumvent a no-go theorem mentioned in [12] on the existence of exact holographic codes with polynomially-decaying correlation functions. Recent work suggests that a bulk-to-boundary code replicating full quantum gravity should be expressed in terms of an approximate error-correcting code [58, 59, 60, 61, 62], and that stabilizer codes in particular only exhibit trivial area operators [63], meaning that quantum superpositions of bulk geometries on subregions are restricted to classical values for the area of the subregion boundaries (Ryu-Takayanagi surfaces). As discussed in [24], this still allows for restricted quantum gravity effects where the bulk subregion accessible from a boundary subregion can fluctuate between configurations with equal area. An interesting implication of our work is that gauge-fixing, which in the holographic picture corresponds to a choice of a given local (quantum) geometry, implies different reconstruction properties and thresholds. For example, the $X$ gauge in the $\{5,4\}$ Evenbly code leads to $A$ tensors describing 4-qubit PME states composed of EPR pairs, resembling the “fixed-area states” described by HaPPY codes [42, 43]. Different gauge choices can thus be seen as a modification of bulk entanglement to suppress or enhance certain types of operators under a bulk RG flow. It is therefore possible that understanding the suitability of holographic codes under different types of errors will also lead to new insights into how quantum gravity emerges in such tensor network models.
IV.2 Error Correction

### v2

@@ 3.3 Evenbly Codes as Subsystem Codes
Let us briefly review subsystem codes. In general quantum error correction codes, one typically divides the physical Hilbert space $\mathcal{H}_{p}$ as $\mathcal{H}_{p}=\mathcal{H}\oplus\mathcal$$\bar${H}}, where $\mathcal{H}$ represents the code subspace, and $\mathcal$$\bar${H}} is the complement. For more general subsystem codes, it is generally taken that $\mathcal{H}$ can be decomposed as $\mathcal{H}=\mathcal{H}_{L}\otimes\mathcal{H}_{G}$ [54], where $\mathcal{H}_{L}$ and $\mathcal{H}_{G}$ represent the logical and gauge subsystems of the code subspace, respectively. In this paradigm, logical information is encoded into $\mathcal{H}_{L}$, while the subsystem $\mathcal{H}_{G}$ is used to help diagnose and correct errors in $\mathcal{H}_{L}$. This aim can be accomplished by defining a gauge group $\mathcal{G}$, of which the total stabilizer group $\mathcal{S}$ is the center of $\mathcal{G}$, i.e. $\{iI,\mathcal{S}\}=\mathcal{Z}(\mathcal{G}):=C(\mathcal{G})\cap\mathcal{G}$. Here $iI$, $\mathcal{Z}(\cdot)$, and $C(\cdot)$ represent the identity operator (up to a phase), the center, and the centralizer, respectively [41]. Up to phase factors, therefore, elements from $\mathcal{G}$ are either stabilizer operators (and act trivially on $\mathcal{H}_{L}\otimes\mathcal{H}_{G}$) or operators which act non-trivially on $\mathcal{H}_{G}$. Lastly, due to the subsystem structure, two types of logical operators arise. Bare logical operators are those which act non-trivially only on $\mathcal{H}_{L}$, while dressed logical operators act non-trivially on $\mathcal{H}_{L}\otimes\mathcal{H}_{G}$ [53, 55].
@@
We show below that this constant-rate version of the Evenbly code yields a rate of $R=k/n=1/8$; for comparison, the trivial gauge subsystem variant (i.e. the maximum number of bulk logical indices are used for storing logical information and gauge group is empty) of the $\{5,4\}$ Evenbly code exhibits an asymptotic rate of $R=0.866$. Such finite encoding rates are a common feature of codes defined on hyperbolic manifolds [56, 8]. Both of the code variants’ threshold properties will be evaluated in Section 3.6. In both cases, $\mathcal{H}_{\text{bulk}}$ refers to all other logical qubits partitioned into the gauge subsystem.
Figure 6: Gauged Evenbly codes considered in this paper. (a) A zero-rate Evenbly code where only the central logical qubit (red dot) is considered, all other bulk legs being gauge-fixed (gray disks). (b) A constant-rate Evenbly code, where each logical qubit occupies a unique crossing point in the bulk. Here, hyperbolic geodesics only pass over sites that are gauge-fixed. An example for such a pair of geodesics for one logical qubit is shaded in blue. (c) In the $X$ gauge, all the gauge-fixed tensors act as pairs of overlapping EPR pairs, mapping each logical qubit to four unique physical boundary sites under the initial $\llbracket 4,1,2\rrbracket$ code. Any other gauge mixes the encodings of logical qubits. (d) A variant of the code in (b), with half of all logical qubits on every layer gauge-fixed. We examine this code’s threshold under erasure in Section 3.6.
@@
3.4 Rate and Distance Scaling
Calculating the asymptotic rate of a $\{5,4\}$ Evenbly code can be done by following the prescriptions found in [14, 8, 47], and is related to an inflation procedure for hyperbolic tilings known as vertex inflation [46, 47]. As in these previous examples, rates can be evaluated by comparing the number of tensors in the bulk with the total number of boundary sites.
Under vertex inflation, one defines each layer of the tiling as a closed loop of vertices (gray disks in our figures), with the initial (zeroth) layer consisting only of the central tensor. One then distinguishes the vertices, each of which hosts a logical qubit, by the number of outgoing planar legs pointing from each vertex directly towards the next layer; at the last layer, these open legs support the physical qubits. We denote vertices with one to four outgoing legs with $\alpha,\beta,\gamma$ and $\delta$. The letter $\delta$ only appears as the central (seed) tensor with four planar legs, and $\gamma$ never appears at all. The sequence of vertices on subsequent layers can be found by applying the substitution rules
@@ 3.4 Rate and Distance Scaling
For example, applying two inflation steps to the central tensor produces the sequences
$\delta\mapsto(\alpha\beta\beta)^{4}\mapsto(\alpha\beta\alpha\beta\beta\alpha\beta\alpha\beta\beta\alpha\beta)^{4}\ ,$ (9)
where we have used powers to indicate repetitions of the sequence. This corresponds to the tilings shown in Figure 1 and Figure 6: We find that the first and second layer around the central vertex indeed contain $12$ and $48$ vertices, respectively, with the latter consisting of $n_{\alpha}=20,n_{\beta}=28$ vertices of each type. It follows that this tensor network has $n=n_{\alpha}+2n_{\beta}=76$ physical boundary sites. More formally, vertex inflation determines a substitution matrix $M$; every column of $M$ represents the number of sites of each type that arise in every subsequent layer as a result of applying the inflation rules. Combining this matrix with the vector $\eta=(n_{\alpha},n_{\beta},n_{\delta})^{\text{T}}$ which represents the number of sites at the $L^{\text{th}}$ layer, we construct a recursion relation of the form
$\eta^{(L+1)}\equiv\begin{pmatrix}n_{\alpha}^{(L+1)}\\ n_{\beta}^{(L+1)}\\ n_{\delta}^{(L+1)}\end{pmatrix}$ $=\begin{pmatrix}1&2&4\\ 1&3&8\\ 0&0&0\end{pmatrix}\begin{pmatrix}n_{\alpha}^{(L)}\\ n_{\beta}^{(L)}\\ n_{\delta}^{(L)}\end{pmatrix}$
$\equiv M\eta^{(L)}\ .$ (10)
@@
This is the rate of the max-rate Evenbly code in which all logical qubits are gauge-free. However, in Section 3.6, we treat another variant, where a number of logical qubits are gauge-fixed. The symmetries of the $\{5,4\}$ tiling allow for a particularly elegant gauge choice: Each vertex in the tiling is an orthogonal crossing point of two geodesics through the hyperbolic disk (shown in Figure 6(c)), with the new geodesics at each layer passing either over a series $\beta-\beta$ of two vertices or over $\beta-\alpha-\beta$ of three. Choosing only every second $\beta$ vertex at each layer as the location of a gauge-free logical qubit then associates each vertex with its “own” pair of geodesics than cross no other gauge-free logical qubit. As each geodesic ends in two physical sites on the boundary, it is unsurprising to find that this code has the constant rate
@@
In our threshold studies we will consider an even sparser version of this model, where only every fourth $\beta$ vertex remains gauge-free (Figure 6(d)), leaving the rate $R^{\prime}=1/8$.
We now consider code distances, i.e., the minimum physical weight of any logical operator, distinguishing between bit distances $d_{\text{bit}}$ of single-qubit logical operators in the bulk basis and word distances $d_{\text{word}}$ of arbitrary multi-qubit logical operators. In most literature, $d_{\text{word}}$ is simply referred to as the distance $d$ of the code [58]. For the Evenbly code, both bit and word distances heavily depend on the choice of gauge: In the max-rate case, acting with two physical Pauli operators can create a logical operation arbitrarily deep in the bulk, leading to $d_{\text{word}}=2$. If the two Pauli operators act on neighboring physical qubits, they will affect only logical qubits adjacent to the boundary (e.g. $Z_{j}Z_{j+1}$ acting as a $\bar{Z}$ on a single tile); however, if they act further apart, the result is a string of logical operators along a shortest path throughout the bulk, a feature that is related to the appearance of non-trivial 2-point functions in such codes [24].
@@
$w^{(L)}(\bar{X})$ $=(1+\sqrt{2})^{L}+(1-\sqrt{2})^{L}\ ,$ (22a)
$w^{(L)}(\bar{Z})$ $=(1+\sqrt{2})^{L+1}+(1-\sqrt{2})^{L+1}\ .$ (22b)
@@
$d_{\text{bit}}\sim(1+\sqrt{2})^{L}$ $=\left(\frac{n}{2+2\sqrt{3}}\right)^{\log_{2+\sqrt{3}}(1+\sqrt{2})}$
@@
The distance for $\bar{Z}$ scales similarly, with an additional prefactor of $1+\sqrt{2}$. However, in general the distance of the entire code is upper-bounded as $d_{\text{word}}\leq d_{\text{bit}}$. To achieve a more useful code, we would like to increase $d_{\text{word}}$ (towards the upper bound dictated by $d_{\text{bit}}$) for a greater resilience against errors so that targeted logical operations are easier to perform.
This condition is exactly what gauge-fixing some logical qubits of the Evenbly code achieves: In the most extreme example of the zero-rate code with only a single gauge-free logical qubit in the center of the tiling, $d_{\text{word}}=d_{\text{bit}}$ while being upper-bounded by the max-rate result from Section 3.4. This upper bound may not be tight, as the availability of additional stabilizer terms on gauge-fixed tiles allows for more compact representations of logical operators on the boundary. Exactly computing these distances is a computationally difficult problem, though we expect that gauge-fixing in a Z or $Y$ gauge will reduce the power coefficient in Section 3.4 somewhat below $0.67$. Note that the arguments for a $d_{\text{bit}}=O(n^{0.67})$ bound should apply to a logical qubit at any bulk location as layers are added, though prefactors will be highly depend on its distance to the boundary; for example, a logical qubit on the last layer (assumed not to be gauge-fixed) will always have weight-$2$ logical operators.
@@
3.5 Erasure Threshold for a Zero-Rate Code
We analyze the threshold performance of the $\{5,4\}$ Evenbly code using two different decoding strategies for erasure errors: A greedy reconstruction method generalizing the approach of [8] adapted to our GHZ-based Evenbly code, in which the bulk logical qubits are iteratively reconstructed from the boundary and whose threshold we denote by $p^{\text{greedy}}_{\text{erasure}}$, and a Gaussian elimination algorithm adapted from [9] with threshold $p^{\text{gaussian}}_{\text{erasure}}$. The implementation and simulation details for both algorithms can be found in Appendix E. To summarize, greedy reconstruction relies on simple geometric steps but is less effective than the Gaussian elimination approach, which considers all possible physical representations of logical operators.
@@ 3.5 Erasure Threshold for a Zero-Rate Code
Simulation results for recovery rates with both decoding algorithms under an erasure channel are shown in Figure 7, considering only the recovery of the central logical qubit. In subfigures 7(a)-(d), the threshold results $p_{\text{erasure}}$ are shown for each choice of gauge fixing discussed in Section 3.3. No gauge fixing (subfigure 7(a)) corresponds to the max-rate case, whereas subfigures 7(b)-(d) consider different gauges of the zero-rate code. In agreement with the discussion of a ququart Evenbly code in [24], the max-rate case is highly susceptible to erasure errors and does not display a threshold.
@@
The logical $Y$ and $Z$ gauges (subfigures 7(c)-(d)) are different: Here we find a threshold close to the maximal value of $50\%$ for full reconstruction, with the simpler greedy reconstruction only achieving around $24\%$ and $40\%$, respectively. This relative underperformance of the greedy algorithm is due to its reconstruction of the bulk algebra on edges tile-by-tile, which may be prevented by erasure errors even though the logical operators of the central qubit are still recoverable. For this reason, the greedy recovery probability always lower-bounds full recovery, as already seen in other work [14, 9]. Conversely, the Gaussian elimination algorithm poses an upper bound to the recovery probability, as it has been shown to be optimal for erasure noise [9, 27].
3.6 Erasure Threshold for a Constant-Rate Code
As shown in Figure 6(b), one may also gauge-fix some of the logical qubits while keeping the rate of the code finite, leading to the constant-rate setting described in Section 3.3. Keeping the two graph geodesics across any logical qubit gauge-fixed on all other bulk sites, we potentially increase the protection of each logical qubit from erasures closest to it on the boundary. In the $X$ gauge, this behavior directly generalizes the zero-rate case: Again, the gauged sites turn into EPR pairs that directly map a $\llbracket 4,1,2\rrbracket$ code onto four unique physical qubits, but with $k=n/4$ logical qubits instead of one. This is visualized in Figure 6(c).
The case of either the $Y$ or $Z$ gauge again leads to non-trivial $L$ dependence and an asymptotic threshold, as in the zero-rate case; the recovery curves are shown in Fig. 8. They can be boosted by adding completely $Y$ gauge-fixed layers of tensors to the constant-rate code, pushing them close to $50\%$ while decreasing the rate accordingly, as we report in Table 1. Note that for such codes, a threshold for the central logical qubit implies a threshold for all the others, as the threshold in a holographic code depends on how much each tensor network layer suppresses errors [9, 28, 13, 8].
@@ 3.6 Erasure Threshold for a Constant-Rate Code
3.7 Threshold for Depolarizing and Pure Pauli Noise
Figure 9: Recovery curves for the zero-rate $\{5,4\}$ Evenbly code for depolarizing and biased Pauli noise, under $Z$ gauge fixing. (a) A threshold appears at approximately $19.1\pm 0.94\%$ for depolarizing noise. (b) For pure Pauli $X$ noise, a preliminary threshold appears at approximately $50\%$. The same behavior was observed for pure Pauli $Y$ and $Z$ noise as well.
@@ 3.7 Threshold for Depolarizing and Pure Pauli Noise
In Figure 9, (a) displays the recovery threshold $p_{\text{depo}}$ obtained as a result of varying the physical noise parameter. As is shown, a threshold appears in the region of approximately $p_{\text{depo}}\approx 19.1\pm 0.94\%$; this threshold value is very competitive with known code-capacity thresholds for the surface code, color code, and the holographic $\llbracket 6,1,3\rrbracket$ code [6, 13, 30]. Subfigure 9(b) displays the results obtained for pure Pauli-X noise using the integer-optimization decoder (we have additionally tested pure Pauli-Y and Pauli-Z noise, and have found the same results). As can be ascertained, preliminary estimates indicate pure Pauli-X noise thresholds of approximately $50\%$, in line with the results obtained for the erasure channel. However, as we were limited in the computational power and time available, we report these findings here as preliminary, with three crossing points. However, if confirmed, it would seem that the Evenbly code is capable of reaching and exceeding known achievable bounds for random zero-rate quantum codes (known as the zero-rate hashing bound [31, 30]).
4 Discussion
4.1 qCFTs & Holography
@@ 4 Discussion
The construction presented in this work is one among many possible realizations of Evenbly codes; indeed, it was shown in [26] that many possible tensor decompositions for the original HTN ansatz exist, even those which are non-unitary in terms of individual tensor properties. However, constructing a quantum error-correcting code with non-unitary tensors may prove to be very challenging, and even more so when trying to determine relevant error correction properties of the resulting code.
@@
Although we have realized an exact holographic code which exhibits genuine polynomially-decaying boundary-to-boundary correlation functions, we have yet to investigate whether CFT data relevant for a known minimal-model CFT [49], or a quasiperiodically disordered qCFT [48], can appear as a result of the gauge-fixing techniques described. Indeed, recent work [61] a MERA-style tensor network was proposed in order to capture certain aspect of discretized AdS/CFT, in addition to recent efforts to parameterize and classify central charges of qCFTs based on their tiling structure [47]. We leave this issue for future work.
@@
Evenbly codes at their core circumvent a no-go theorem mentioned in [12] on the existence of exact holographic codes with polynomially-decaying correlation functions. Recent work suggests that a bulk-to-boundary code replicating full quantum gravity should be expressed in terms of an approximate error-correcting code [62, 63, 64, 65, 66], and that stabilizer codes in particular only exhibit trivial area operators [67], meaning that quantum superpositions of bulk geometries on subregions are restricted to classical values for the area of the subregion boundaries (Ryu-Takayanagi surfaces). As discussed in [24], this still allows for restricted quantum gravity effects where the bulk subregion accessible from a boundary subregion can fluctuate between configurations with equal area.
An interesting implication of our work is that gauge-fixing, which in the holographic picture corresponds to a choice of a given local (quantum) geometry, implies different reconstruction properties and thresholds. For example, the $X$ gauge in the $\{5,4\}$ Evenbly code leads to $A$ tensors describing 4-qubit PME states composed of EPR pairs, resembling the “fixed-area states” described by HaPPY codes [44, 45]. Different gauge choices can thus be seen as a modification of bulk entanglement to suppress or enhance certain types of operators under a bulk RG flow. It is therefore possible that understanding the suitability of holographic codes under different types of errors will also lead to new insights into how quantum gravity emerges in such tensor network models.
4.2 Error Correction